OpenStudy (anonymous):

explain how to Differentiate f(x)=-5x(2X-3)^4

OpenStudy (anonymous):

the form is off...

OpenStudy (anonymous):

are you asking -5x * 2 * -3^4

OpenStudy (anonymous):

if so... -3^4 is 81... times 2 is 162.. times -5 is -810

OpenStudy (anonymous):

im asking to find the derivative of\[-5x(2x-3)^{4}\]

OpenStudy (anonymous):

but I doubt that is what your looking for...

OpenStudy (anonymous):

product rule with the differentiation of one being the chain rule

OpenStudy (anonymous):

chain rule... exactly

OpenStudy (anonymous):

a'b + b'a

OpenStudy (anonymous):

\[-5x*(2x-3)^4\]

OpenStudy (anonymous):

so start out by dividing the two up... -5x and (2x-3)^4

OpenStudy (anonymous):

call them a and b

OpenStudy (anonymous):

u = -5x v = (2x-3)^4 v' = 8(2x-3)^3 u' = -5 the chain/product rule is: uv'+vu' just attached the givens, :))

OpenStudy (anonymous):

kumar is correct..

OpenStudy (anonymous):

multiply the u and v', then the u' and v, then add the two...

OpenStudy (anonymous):

haha i'm afraid to foil the parenthesis lmfao

OpenStudy (anonymous):

that would be ugly...

OpenStudy (anonymous):

Dannibe, do you understand?

OpenStudy (anonymous):

oh yes it would

OpenStudy (anonymous):

no i can't get the right answer

OpenStudy (anonymous):

=]

OpenStudy (anonymous):

actually the 4 wouldn't be cubed

OpenStudy (anonymous):

the answer is supposed to be \[-5(2x-3)^{3}(10x-3)\]

OpenStudy (anonymous):

so your starting with -5(2x-3)^4 - 40x(2x-3)^3 ?

OpenStudy (anonymous):

i dont see this equation simplifying to that.. but let me check

OpenStudy (anonymous):

Thats where I got to. how would you work it after that point?

OpenStudy (anonymous):

ahhhh

OpenStudy (anonymous):

devide the whole line by (2x-3)^3

OpenStudy (anonymous):

pull that to the front..

OpenStudy (anonymous):

the you have (2x-3)^3 * (-5 - 40x*(2x-3))

OpenStudy (anonymous):

distribute the 40x*(2x-3)

OpenStudy (anonymous):

err.... start with -5(2x-3)^4 - 40x(2x-3)^3

OpenStudy (anonymous):

then go to -5(2x-3)^3 * (1-8x)

OpenStudy (anonymous):

look like I fooked it up... reworking

OpenStudy (anonymous):

ok.. the anwser is correct..

OpenStudy (anonymous):

here is how you get there...

OpenStudy (anonymous):

start with -5(2x-3)^4 - 40x(2x-3)^3 factor out -5(2x-3)^3 then you get -5(2x-3)^3 * (2x - 3 + 8x)

OpenStudy (anonymous):

hence your anwser